# A Multi-Perspective Analysis of Memorization in Large Language Models

### v1

@@ 4.4.1 Embedding Dynamics
In this section, we discuss the embedding of generated tokens for sentences memorized to different extents. We collect the hidden state at the last layer of each token. We compute the pair-wise Euclidean distance and Cosine similarity between sentences memorized at different extents across each generation step. Then, we perform dimension reduction on the embedding and visualize it in the 2D space using PCA.
Figure 5: Embedding Dynamics Across Different Model Size. Score $x$ means $x$ tokens are memorized.
@@
  • The cosine similarity remains relatively stable across different steps between sentences memorized at different extents. This shows that different memorized sentences have different parts in the embedding space and remain relatively stable.
  • Meanwhile, we can see the Euclidean distance decreases with the generation of tokens. Combing with the previous observation that sentences memorized with different extend have a stable Cosine Similarity, we can conclude that with the generation of tokens, the angle between sentence vectors at high dimension space remains stable while the cosine distance decreases, which suggests the sentence vectors is shortened while the angle remains stable.
  • Additionally, we can see that the large-size model has a higher Euclidean distance and lower Consine Similarity. The high Euclidean distance is caused by the expansion of hidden size (512 for 70m, 2048 for 1b, etc.). As for the decrease of low Consine Similarity, this can also be attributed to the hidden size as the increase in hidden size expands the embedding space. This gives more expressivity to the embedding space, as the Euclidean distance between sentence embedding also expands, which leads to the decrease of Cosine Similarity. This may also help to explain the performance gap between different model sizes. Large-size models have more expressivity in the embedding space; thus, different sentences with different meanings are more clearly distributed with fewer neighbors in the space. The smaller embedding space in a small model makes the mix of embedding easier, leading to ambiguity in the embedding space.
@@ 4.4.2 Generation Dynamics and Entropy
In this section, we discuss how the model generates the memorized and unmemorized sentences, e.g., the generation dynamics. This means that the model exhibits certain behaviors when it generates memorized sentences or unmemorized sentences. Or does the LLM generate memorized and unmemorized sentences in the same way? To discuss this problem, we collect the generation probability at each token when the LLM generates memorized, unmemorized, and half-memorized sentences, calculate the entropy across 10,000 sentences, and plot the entropy at each token in Figure.
@@
From the Figure 6 we can tell:
  • First, we can see that the entropy is different depending on the extent of memorization. We can clearly see that an unmemorized sentence has a higher average entropy at each token than an unmemorized sentence. This shows that when the LLM generates unmemorized sentences, they are more unconfident about generating tokens.
  • Additionally, we can tell that when the model starts to generate unmemorized tokens, we can see a significant increase in entropy. This boundary effect is very obvious. This suggests that when the model generates unmemorized sentences, their entropy will suddenly increase instead of decrease.
  • We can also notice that the boundary for generating unmemorized sentences is mostly consistent, which means we did not see multiple boundaries. This suggests that even for half-memorized sentences, the model tends to generate continuous memorized tokens then, followed by unmemorized tokens instead of generating mixed memorized and unmemorized tokens. This suggests that the LLM memorizes sentences at a chunk level rather than a word level.
  • We can also see that the entropy will decrease with the increase in model size. This suggests that the larger size model is more confident about the generation of tokens compared to the small size model.
@@ 4.5 Prediction of Memorization
In this section, we train a Transformer model to predict the memorization of tokens from the model’s embedding and statistic features and analyze the results. We discuss whether Transformer is capable of this task and the difficulty of predicting the memorized token at different model sizes as in the real world, similar context sentences can also trigger memorized texts Stoehr et al. (2024) and the generated token can also be a paraphrase of the actual continuation Ippolito et al. (2023).
@@ 4.5.1 Results on Memorization of Prediction
The training is formalized by receiving both the last layer’s embeddings at each step and statistics (e.g., the entropy) then the Transformer is required to generate a binary classification label at each index, predicting whether the token at this index is memorized. The results are presented in Table 2. From the table, we can obtain the following results:
  • First, regarding Token-level accuracy, we can see with a naive Transformer model, the token-level accuracy can reach $80\%$ and even higher, showing that the memorization label of $80\%$ tokens can be correctly classified. The Full-level accuracy is relatively low as this requires a correct prediction at every token. However, the high token level accuracy at such a naive model shows the potential that prediction of memorization at the decoding step is possible.
  • Secondly, we showed that with the increase of the model size of LLM, the Token-level accuracy and Full-level accuracy increase rather than drop. This suggests that it is easier to predict the memorization behavior of larger-size models rather than small-size models.
  • Thirdly, we found that with the increase in continuation size, the Token level accuracy gradually increases, possibly due to the increase in training data. For example, if the length to be predicted is $64$, the token used to train the model is four times higher than the prediction length of $16$. However, we see a drop in the Full accuracy as the increase in prediction length leads to more tokens to be predicted; thus, reaching Full accuracy is harder in a longer prediction length.
@@ 4.5.2 Analysis of Full Accuracy
  • For any model size, most sentences that are fully predicted share low memorization scores. Especially, the 00 score (unmemorized) sentence occupies the largest portion among fully predicted sentences. Meanwhile, we can see that sentences with high memorization scores are hard to be fully predicted. We can see that the $1$ score (memorized) sentence has the lowest share among fully predicted sentences, showing that memorized sentences are hard to predict.
@@
  • From previous Figure 6, for the memorized sentence, the entropy decrease after decoding is less significant than for the unmemorized sentence, whose entropy will suddenly increase. Additionally, the decrease in entropy is less significant with the increase in model size as the large model is more confident about the decoding, leaving less space for the entropy decrease. Since entropy is part of the feature we input to the model, the aforementioned entropy decreases, making a prediction of memorized sentences easier. The less significant decrease of entropy in memorized sentences helps to explain why the prediction of memorized sentence
@@ 4.6 Conclusion
In this research, we comprehensively discussed the memorization behavior of LLMs across different perspectives. From a general level, we conclude statistics regarding the memorization behavior of LLMs, which not only aligns with previous research but also extends to sentences with lower memorization scores. From the embedding level, we obtained the embedding dynamics of sentences with different memorization scores. We found that the cosine similarity stays stable among sentences with different memorization scores, but the Euclidean distance decreases. The increase in model size leads to a decrease in cosine similarity and an increase in Euclidean distance. At the decoding dynamics, we conducted an analysis at the distribution level and calculated the entropy. We found that when the LLM starts to generate an unmemorized sentence, the entropy will suddenly rise at the boundary between memorized and unmemorized parts. On the contrary, when the model generates memorized sentences, the entropy will also have a sudden drop, though not obvious, comparing the generation of unmemorized parts. Finally, we trained a Transformer model to predict the memorization. We found that, Through a detailed case study, we analyzed the prediction behavior and showed that context-based memorization prediction is possible.

### v2

@@ 4.4.1 Embedding Dynamics
In this section, we discuss the embedding of generated tokens for sentences memorized at different extents. We collect the hidden state at the last layer of each token. We compute the pair-wise Euclidean distance and Cosine similarity between sentences memorized at different extents across each generation step. Then, we perform dimension reduction on the embedding and visualize it in the 2D space using PCA.
Figure 5: Embedding Dynamics of Across Different Model Size. Memorized Token Num $x$ means $x$ tokens generated align with the true continuation.
@@
  • The cosine similarity remains relatively stable across different encoding steps between sentences memorized at different extents. Meanwhile, from the movement of those points, we can see the Euclidean distance decreases with the generation of tokens. Combined with the previous observation that sentences memorized with different extend have a stable cosine similarity, we can conclude that with the generation of tokens, the angle between sentence vectors at high dimension space remains stable while the cosine similarity decreases.
  • Additionally, we can see that for highly memorized sentences, even when they are not totally memorized. They are still close in the embedding space indicating that the generated content may be both semantic and lexically similar to the memorized content.
  • Additionally, we can see that the large-size model has a higher Euclidean distance and lower Consine Similarity. The high Euclidean distance is caused by the expansion of hidden size (512 for 70m, 2048 for 1b, etc.). As for the decrease of the cosine similarity, this can also be attributed to the hidden size as the increase in hidden size expands the embedding space. This gives more expressivity to the embedding space, as the Euclidean distance between sentence embedding also expands, which leads to the decrease of Cosine Similarity. This may also help to explain the performance gap between different model sizes. Large-size models have more expressivity in the embedding space; thus, different sentences with different meanings are more clearly distributed with fewer neighbors in the space. The smaller embedding space in a small model makes the mix of embedding easier, leading to ambiguity in the embedding space, thus degrading the performance.
@@ 4.4.2 Generation Dynamics and Entropy
In this section, we discuss how the model generates the memorized sentences with different memorization scores, e.g., the generation dynamics. As we have found the boundary effect in the input level, we also want to discuss the question of does the model exhibit certain behaviors when it generates memorized sentences or unmemorized sentences? To discuss this problem, we collect the generation probability at each token when the LLM generates memorized, unmemorized, and half-memorized sentences, calculate the entropy across 10,000 sentences, and plot the entropy at each token in Figure 6 and we can obtain the following from it:
  • First, we can see that the entropy is different depending on the extent of memorization. We can clearly see that an unmemorized sentence has a higher average entropy at each token than an unmemorized sentence. This shows that when the LLM generates unmemorized sentences, they are more unconfident about generating them.
  • Additionally, when the model starts to generate unmemorized tokens, we can see a significant increase in entropy, showing such boundary effect also exists in the decoding steps while behaving oppositely compared to the boundary effect in frequency analysis. For the memorized sequences, we can also see a drop in entropy when decoding them. This corresponds to the whole memorized tokens and the first half of the half-memorized sequences; the entropy drops around the index of 32, which is the first generated token. However, we can see the drop of entropy for the memorized sentence is not samely significant as the half-memorized sentence, showing that even for the memorized content, the model is more confident when the sentence is fully memorized, leading to lower entropy.
  • We can also notice there is only one boundary at unmemorized or memorized sentences, further confirming that the model tends to memorize tokens continuously.
  • We can also see the entropy decrease with the increase in model size. This suggests that the larger size model is more confident about the generation of tokens compared to the small size model. Additionally, we can even see that the entropy is also lower in the context of memorized tokens, showing that the entropy of context also relates to how a sentence is memorized.
@@ 4.5 Prediction of Memorization
In this section, we train a Transformer model to predict the memorization of tokens from the LLM’s embedding and statistic features and analyze the results. We discuss whether Transformer is capable of this task and the difficulty of predicting the memorized token at different model sizes. As in the real world and indicated by our experiment results in previous sections, similar context sentences can also trigger memorized texts Stoehr et al. (2024), and the generated token can also be a paraphrase of the actual continuation Ippolito et al. (2023).
@@ 4.5.1 Results on Memorization of Prediction
The training is formalized by receiving both the last layer’s embeddings at each step and statistics (e.g., the entropy). Then, the Transformer is required to generate a binary classification label at each index, predicting whether the token at this index is memorized. The results are presented in Table 2. From the table, we can obtain the following results:
  • First, regarding Token-level accuracy, we can see with a naive Transformer model, the token-level accuracy can reach $80\%$ and even higher, showing that the prediction of memorization at the token level is easy. The Full-level accuracy is low as this requires a correct prediction at every token.
  • Secondly, we showed that with the increase of the model size of LLM, the Token-level accuracy and Full-level accuracy increase rather than drop. This suggests that it is easier to predict the memorization behavior of larger-size models rather than small-size models as the embedding distance of large models is further, thus making the classification easier.
  • Thirdly, we found that with the increase in continuation size, the Token level accuracy gradually increases, possibly due to the increase in training data. For example, if the length to be predicted is $64$, the token used to train the model is four times higher than the prediction length of $16$. However, we see a drop in the Full accuracy as the increase in prediction length leads to more tokens being predicted. This makes reaching Full accuracy harder as the tokens required predicted grow.
@@ 4.5.2 Analysis of Full Accuracy
  • For any model size, most sentences that are fully predicted share low memorization scores. Especially, the 00 score (unmemorized) sentence occupies the largest portion among fully predicted sentences. Meanwhile, we can see that sentences with high memorization scores are hard to fully predict. We can see that the $1$ score (memorized) sentence has the lowest share among fully predicted sentences, showing that memorized sentences are hard to predict.
@@
  • The reason why a sentence with a low memorization score can be easily predicted can be due to the difference of significance at the boundary. As obtained from our previous experiments, the boundary effect of unmemorized sequences is more significant compared to memorized sequences in either token frequency or entropy of the generated token. Especially we can see that the boundary effect of memorized sequences decreases with model size while increasing for unmemorized sequences in previous results. This helps to explain why predicting memorization is easier in small-size models but harder in large-size models because the boundary effect becomes insignificant when we increase the model size. Similarly, the increase of boundary effect for unmemorized sentences with model size makes them easier to predict.
@@ 4.6 Conclusion
In this research, we comprehensively discussed the memorization behavior of LLMs across different perspectives. From a statistical level, we analyzed the memorization behavior of LLMs and expanded the analysis to sentences with lower memorization scores. Further transition analysis revealed the transition dynamics when trained in a large model, showing that most sentences stay in the original state while the memorized state is hard to transition to the unmemorized state. In the input level, through frequency analysis, we found the positive and negative boundary effect when generating memorized and unmemorized tokens and showed the LLM tends to memorize longer sequences rather than short pieces. In the output decoding level, in the embedding level, we found sentences with different memorization scores clusters in the embedding space, and the close distance of sentences with high memorization scores indicates the existence of paraphrase memorization. At the decoding dynamics, we conducted an analysis at the distribution level and calculated the entropy. We found that when the LLM starts to generate an unmemorized sentence, the entropy boundary effect also exists and significantly changes with the model size and the memorization score of the corresponding sentence. Finally, we trained a Transformer model to predict the memorization. We found that prediction at the token level is easy, while prediction at the sentence level is challenging. Through analysis of correctly fully predicted samples, we found unmemorized tokens are easy to predict while memorized tokens are hard to predict, and such a phenomenon can be explained by the significance of the boundary effect.